# The TRGB--SBF Project. III. Refining the HST Surface Brightness Fluctuation Distance Scale Calibration with JWST

### v1

@@ 3.1 Updated Calibration
Figure 2: Absolute fluctuation magnitude $\overline{M}_{110}$ as a function of DECam galaxy color converted to the PanSTARRS photometric system for the eight galaxies with TRGB distances (Anand et al., 2024b, 2025) plus six additional galaxies in the Fornax and Virgo clusters previously used to calibrate F110W SBF (blue points). The fit shown was performed using the “emcee” package in Python, iteratively accounting for uncertainties in both axes. The blue line is the best fit to the blue points; the blue envelopes mark the 1-$\sigma$ and 2-$\sigma$ uncertainties in slope and intercept. The absolute magnitudes calculated using the individual TRGB distance moduli are also plotted for five of the galaxies with PanSTARRS colors (green points) and eight galaxies with DECam colors (black points) to show that slopes derived from the individual and cluster distances are consistent.
3.2 Updated SBF Distances
@@
We are confident that the result of applying the new calibration based on the JWST TRGB measurements, including improved optical color measurements, does not change the conclusions of Blakeslee et al. (2021) regarding the value of $H_{0}$ in any significant way. The resulting value of $H_{0}$ using the updated zero point and the published calibration slope gives $73.81$ km s-1 Mpc-1 instead of $73.44\pm 0.7$ km s-1 Mpc-1, the value found in our 2021 work when using only the Cepheid calibration. Since we are only changing the calibration and not the input $\overline{m}_{110}$ fluctuation magnitudes or velocities, the statistical error in the value of $H_{0}$ is unchanged. A similar 0.5% increase would result for the other values of $H_{0}$ calculated from IR SBF distances (Garnavich et al., 2023; Blakeslee et al., 2021).
@@ 3.3 Systematic Uncertainties
The SBF distance scale is primarily limited by systematic uncertainties. The Cepheid calibration of SBF injected a 0.09 mag systematic uncertainty (Blakeslee et al., 2021) that we remove here and replace with that of the NGC 4258 megamaser and JWST TRGB calibration. The systematic uncertainty associated with the maser distance to NGC 4258 is 1.48% (0.032 mag, Reid et al. 2019). The TRGB measurement uncertainty of 0.033 mag for NGC 4258 from Anand et al. (2024a) acts as an additional systematic error in the TRGB distances of other galaxies that are tied to it. This source of systematic error cannot be reduced further without observations of additional fields around NGC 4258 or other geometrical anchors besides NGC 4258. The link from TRGB to SBF contains the zero point uncertainty of 0.016 mag from Equation 2 above. Combined in quadrature, these add up to 0.049 mag. The intrinsic scatter in the properties of the TRGB feature (and how it is measured) has not been fully quantified. Anand et al. (2024a) estimated a scatter of 0.02 mag in their study, and estimates from Riess et al. (2024) for a sample of spiral galaxies observed in multiple filters with JWST suggest larger values from 0.01 to 0.08 mag. The elliptical galaxies we observed span a smaller range in age and metallicity and should have modest intrinsic scatter; we conservatively include another 0.04 mag in our error budget.
Taken together, these terms give a total systematic uncertainty on the TRGB calibration of HST SBF of 0.063 mag, or 2.9% in distance. Combining this with the estimated 1% uncertainty in the overall velocity flow of the volume covered by our sample (see Blakeslee et al., 2021), the total systematic uncertainty in $H_{0}$ becomes 3.1%, or 2.3 km s-1 Mpc-1 in $H_{0}$. This is a significant improvement over the previous Cepheid-based calibration of 0.09 mag and will improve further with additional geometrical anchors tied to TRGB with Gaia and with more links between JWST TRGB and HST SBF.
@@ 4 Conclusions
This updated calibration is a first step in establishing a new distance ladder with JWST that will be independent of the traditional rungs that use the LMC distance, Cepheids and type Ia supernovae. We are optimistic that additional galaxies will be observed by both JWST and HST, increasing the overlap between the two telescopes both for TRGB and SBF measurements, and further reducing systematic uncertainties.
JWST imaging observations of 39 targets in the Coma cluster (likely including  ${\sim\,}50$ or more elliptical galaxies) have been scheduled for Cycle 3 (GO-5989; Jensen et al. 2024) and will provide a robust determination of the calibration slope in a cluster that is foundational to the distance scale. Additional geometrical anchors securing the absolute magnitude of the TRGB will further enhance the precision of the SBF method. The combination of HST and JWST SBF measurements will create a new distance ladder based solely on old, metal-rich populations in early-type galaxies and free from many of the uncertainties in the Cepheid+SN distance ladder.
Acknowledgments. This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. The results presented here used data associated with HST programs GO-11711, 11712, 12450, 14219, 14654, 15265, 14771, 14804, and 15329. J. Jensen and M. Cowles acknowledge partial support from STScI under grants HST-GO-16262 and JWST-GO-03055. This work also used observations made with the NASA/ESA/CSA James Webb Space Telescope program GO-3055. J. Blakeslee is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation. M. Cowles acknowledges support from The Barry Goldwater Scholarship and Excellence in Education Foundation and the National Science Foundation. M. Cantiello acknowledges support from the project “INAF Exploration of Diffuse Galaxies with Euclid” (INAF-Large Grant 12-2022, P.I. Leslie K. Hunt), and by Italian Ministry for Education University and Research (MIUR) grant PRIN 2022 2022383WFT “SUNRISE,” CUP C53D23000850006 (P.I. Enrica Iodice). G. S. Anand acknowledges financial support from JWST GO–3055.

### v2

@@ 3.1 Updated Calibration
Figure 2: Absolute fluctuation magnitude $\overline{M}_{110}$ as a function of the extinction-corrected DECam galaxy color $(g{-}z)$0 converted to the PanSTARRS photometric system for the eight galaxies with TRGB distances (Anand et al., 2024b, 2025) plus six additional galaxies in the Fornax and Virgo clusters previously used to calibrate F110W SBF (blue points). The fit shown was performed using the “emcee” package in Python (Foreman-Mackey et al., 2013) that we used to iteratively account for uncertainties in both axes. The blue line is the best fit to the blue points; the blue envelopes mark the 1-$\sigma$ and 2-$\sigma$ uncertainties in slope and intercept. The absolute magnitudes calculated using the individual TRGB distance moduli are also plotted for five of the galaxies with PanSTARRS colors (green points) and eight galaxies with DECam colors (black points) to show that slopes derived from the individual and cluster distances are consistent.
3.2 Updated SBF Distances and $H_{0}$
@@
To determine the updated value of $H_{0}$, we combined the revised SBF distances described above with the CMB-frame group/cluster velocities published by Jensen et al. (2021) without refitting for $H_{0}$. Specifically, we took the mean change in distance computed from 61 galaxies, using the new TRGB zero point and including the updated $(g{-}z)$ colors, and applied it to the value of $H_{0}$ from Blakeslee et al. (2021). The resulting value of $H_{0}$ using the updated zero point and the published calibration slope gives $73.81\pm 0.7$ km s-1 Mpc-1 instead of $73.44\pm 0.7$ km s-1 Mpc-1. Since we are only changing the calibration and not the input $\overline{m}_{110}$ fluctuation magnitudes or velocities, the statistical error in the value of $H_{0}$ is unchanged. The updated Hubble diagram is shown in Fig. 3. The Hubble velocities used to compute this value of $H_{0}$ and for Figure 3 are in the CMB frame and assigned to groups or clusters by Tully (2015).
A similar 0.5% increase would result for the other values of $H_{0}$ calculated from IR SBF distances or velocities (Blakeslee et al., 2021; Garnavich et al., 2023). For example, Blakeslee et al. (2021) calculated a value of $H_{0}=73.8$ km s-1 Mpc-1 using the 2M++ redshifts and peculiar velocity field derived from the density model calculated by Carrick et al. (2015); the 2M++ value of $H_{0}$ using the new zero point is $74.2$ km s-1 Mpc-1.
@@
Figure 3: Hubble diagram using distances derived from the updated SBF zero point calibration from the TRGB measurements and improved optical color measurements. Velocities and SBF apparent magnitudes are from Jensen et al. (2021). The solid symbols are those for which there is no evidence of dust or other sprial structure in the SBF analysis region. This plot is directly comparable to the left panel of Fig. 1 from Blakeslee et al. (2021).
@@ 3.3 Systematic Uncertainties
The SBF distance scale is primarily limited by systematic uncertainties. The Cepheid calibration of SBF injected a 0.09 mag systematic uncertainty (Blakeslee et al., 2021) that we remove here and replace with that of the NGC 4258 megamaser and JWST TRGB calibration. The systematic uncertainty associated with the maser distance to NGC 4258 is 1.48% (0.032 mag, Reid et al. 2019). The TRGB measurement uncertainty of 0.033 mag for NGC 4258 from Anand et al. (2024a) acts as an additional systematic error in the TRGB distances of other galaxies that are tied to it. This source of systematic error cannot be reduced further without observations of additional fields around NGC 4258 or other geometrical anchors besides NGC 4258. The link from TRGB to SBF contains the zero point uncertainty of 0.016 mag from Equation 2 above. Combined in quadrature, these add up to 0.049 mag. The intrinsic scatter in the properties of the TRGB feature (and how it is measured) has not been fully quantified. Anand et al. (2024a) estimated a scatter of 0.02 mag in their study, and estimates from Riess et al. (2024b) for a sample of spiral galaxies observed in multiple filters with JWST suggest larger values from 0.01 to 0.08 mag. The elliptical galaxies we observed span a smaller range in age and metallicity and should have modest intrinsic scatter; we conservatively include another 0.04 mag in our error budget.
Taken together, these terms give a total systematic uncertainty on the TRGB calibration of HST SBF of 0.063 mag, or 2.9% in distance. Combining this with the estimated 1% uncertainty in the overall velocity flow of the volume covered by our sample (see Blakeslee et al., 2021), the total systematic uncertainty in $H_{0}$ becomes 3.1%, or 2.3 km s-1 Mpc-1 in $H_{0}$. While this is only slightly better than the systematic uncertainty reported by Blakeslee et al. (2021), it provides strong confirmation that systematic errors in the SBF distance scale zero point are accurately measured, given that the sources of systematic uncertainty in the two measurements are very different (i.e., Cepheids vs. TRGB). The new TRGB-SBF distance calibration will improve further with additional geometrical anchors tied to TRGB with Gaia and with more links between JWST TRGB and HST SBF.
4 Distance to the Coma Cluster
[…]
@@ 5 Conclusions
This updated calibration is a first step in establishing a new distance ladder with JWST that will be independent of the traditional rungs that use the LMC distance, Cepheids and SNe Ia. We are optimistic that additional galaxies will be observed by both JWST and HST, increasing the overlap between the two telescopes both for TRGB and SBF measurements, and further reducing systematic uncertainties.
JWST imaging observations of 39 targets in the Coma cluster (likely including  ${\sim\,}50$ or more elliptical galaxies) have been scheduled for Cycle 3 (GO-5989; Jensen et al. 2024) and will provide a robust determination of the calibration slope in a cluster that is foundational to the distance scale. Additional geometrical anchors securing the absolute magnitude of the TRGB will further enhance the precision of the SBF method. The combination of HST and JWST SBF measurements will create a new distance ladder based solely on old, metal-rich populations in early-type galaxies and free from many of the uncertainties in the Cepheid+ SNe Ia distance ladder the most important ones of which are crowding (e.g., Riess et al., 2023), Cepheid metallicity effects (e.g., Madore & Freedman, 2025; Bhardwaj et al., 2024), and the effects of dust extinction on both SNe Ia and Cepheids (e.g., Brout & Riess, 2023).
Reducing the systematic uncertainty in the TRGB-SBF distance scale will require additional geometrical anchors besides NGC 4258, particularly using Gaia (Anand et al., 2025; Gaia Collaboration et al., 2016) to precisely measure RGB, Horizontal Branch, and RR Lyrae stellar distances where the TRGB can also be measured. Improvements from an expanded SBF sample are also forthcoming. Three HST programs have recently been completed and add more than 150 new IR SBF measurements. JWST observations of the Coma cluster (Jensen et al., 2024) will soon establish a foundation for the SBF calibration that will extend to much larger distances than is possible with HST, and 11 additional TRGB targets have been approved for the next cycle (Tully et al., 2025). The updated SBF calibration will be applied to JWST observations of galaxies reaching 250 Mpc in cycle 4 (GO-7113). These projects will further reduce the random and systematic uncertainties on $H_{0}$ measured using the extensive HST WFC3/IR SBF dataset and new observations with JWST.
Acknowledgments. This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. The results presented here used data associated with HST programs GO-11711, 11712, 12450, 14219, 14654, 15265, 14771, 14804, and 15329. J. Jensen and M. Cowles acknowledge partial support from STScI under grants HST-GO-16262 and JWST-GO-03055. This work also used observations made with the NASA/ESA/CSA James Webb Space Telescope program GO-3055. J. Blakeslee is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation. M. Cowles acknowledges support from The Barry Goldwater Scholarship and Excellence in Education Foundation and the National Science Foundation. M. Cantiello acknowledges support from the ASI-INAF agreement “Scientific Activity for the Euclid Mission” (n.2024-10-HH.0; WP8420) and from the INAF “Astrofisica Fondamentale” GO/GTO-grant 2024 (P.I. M. Cantiello). G. S. Anand acknowledges financial support from JWST GO–3055.